OpenStudy (anonymous):

help me Solve (7x + 1) – 3x = 7x + (1 – 3x) ? please

hero (hero):

Pretty easy

OpenStudy (anonymous):

if you have to get everythhing out of perentse,"(" ")" and combine liek terms like this (7x+1)-3x=7x+(1-3x) -21x-3x=7x-21x

hero (hero):

n2n3help, no, that's not it

OpenStudy (anonymous):

well can you help ?

hero (hero):

You can either remove the parentheses or use associative property to figure this out.

OpenStudy (anonymous):

im still lost

hero (hero):

If you remove the parentheses, you'll have 7x + 1 - 3x = 7x + 1 - 3x If you use associative property, then: (7x + 1) - 3x = (7x + 1) - 3x Either way, it is pretty obvious to see, when you place like terms on the same side: (7x + 1) - (7x + 1) = 3x - 3x 0 = 0

OpenStudy (anonymous):

so its no solution ?

OpenStudy (anonymous):

@StudentAthlete i think you got it...

hero (hero):

It's not "no solution", its "all solutions"

hero (hero):

Every value of x will work

hero (hero):

Why? Because both sides are equal

OpenStudy (anonymous):

so no solution ?

hero (hero):

Face palm

OpenStudy (anonymous):

lol

hero (hero):

I just told you what the answer is, and you're still confused.

OpenStudy (anonymous):

no cause my answers say x = 1/4 x = 4 This is an identity and is true for all real numbers. There is no solution.

OpenStudy (anonymous):

its out of those four

hero (hero):

You should look up "no solution" in the math dictionary. No solution is what results when no value of x works. In this case, ALL values of x will work for the identity.